OpenStudy (anonymous):

question: write the complex number in standard form square root of -16

OpenStudy (anonymous):



OpenStudy (anonymous):

would it be the same for square root of -121?

OpenStudy (anonymous):

\[\sqrt{a*b}=\sqrt{a}*\sqrt{b}\]

OpenStudy (zzr0ck3r):

sqrt(-16) = sqrt(16)sqrt(-1) = 4i

OpenStudy (anonymous):

the answer is you factor the -1 out of the equation: \[\sqrt{-16}=\sqrt{-1}*\sqrt{16}\]

OpenStudy (anonymous):

if i=\[\sqrt{-1}\] then the final answer is 4i

OpenStudy (anonymous):

can you draw it step by step

OpenStudy (zzr0ck3r):

this is not something you draw sqr(ab) = sqrt(a)*sqrt(b) so the steps are sqrt(-16) = sqrt(16)sqrt(-1) = 4i

OpenStudy (anonymous):